OpenStudy (anonymous):

how do you solve by step for sin(pi/3) + cos(pi/6)

OpenStudy (raden):

memorize, sin(pi/3) = sin60 degrees = 1/2 sqrt(3) cos(pi/6) = cos30 degreess = 1/2 sqrt(3) now, add them

OpenStudy (anonymous):

wait why is it 1/2 √3?

OpenStudy (anonymous):

isnt sin pi/3 = √3/2?

OpenStudy (anonymous):

wait nevermind I GOt IT

OpenStudy (raden):

lol, that's similar ^

OpenStudy (ajprincess):

 It may be helpful to u if u remember these:)